# Disentangling Core and Edge Mechanisms of the Density Limit in DIII-D Negative Triangularity Plasmas

### v3

@@ III.3.3 Clamping of the Edge Density
The preceding evidence indicates that the HFS radiative instability is the cause of the edge density saturation. To test this hypothesis quantitatively, we correlate the local edge density with the instability’s strength, again using the RMS amplitude of the line-integrated HFS density fluctuations () as a proxy.
Figure 9 plots the edge density (at ) against on a log-log scale. The data, compiled from multiple discharges, reveals a dramatic bifurcation in the plasma behavior, defined by two distinct operational regimes that are distinguished by the energy confinement factor, .
@@
This linear relationship breaks down abruptly once the turbulence level surpasses a critical threshold, at which point the system transitions to a degraded-confinement state (low , purple points). While the instability amplitude continues to grow by nearly an order of magnitude, the edge density saturates almost completely, exhibiting a nearly flat scaling ().
This bifurcation provides strong evidence that the edge density limit is not a gradual process, but a sharp transition triggered by the radiative instability. Once initiated, the instability clamps the edge density at a value consistently below the Greenwald limit, leading to a significant increase in HFS turbulence. The apparent overall relationship across both regimes () also masks the sharpness of this critical transition.
@@ III.3.4 Breakdown of Zonal Flow Regulation
The onset of the HFS radiative instability has global consequences, leading to a failure of turbulence self-regulation across the plasma edge. We use BES measurements on the low-field side to observe this breakdown in detail (Fig. 10), revealing the microphysical mechanism responsible for enhanced transport.
The trigger for this breakdown is the transition to the non-adiabatic regime. As the edge cools with the onset of the radiation front ( ms), the dimensionless adiabaticity parameter, , drops below unity [Fig. 10(a)]. This transition is known to fundamentally alter drift-wave turbulence, weakening the mechanisms that generate stabilizing zonal flows [37, 13].
@@
This breakdown in zonal flow regulation directly drives particle loss. The outward turbulent particle flux [, Fig. 10(e)] grows significantly, while the turbulent spreading flux [, Fig. 10(f)] indicates that turbulence intensity is propagating inward from the separatrix. These measurements suggest a compelling narrative: edge cooling initiates a transition to the non-adiabatic turbulent regime, which disables zonal flow regulation and unleashes enhanced transport.
III.4 The Core Density Limit: A Transition to Avalanche-like Turbulence
While the edge density saturates below the Greenwald limit, the core density continues to rise, often reaching much higher values than . This decoupling implies that a separate, core-localized mechanism must govern the core density profile. Here, we present evidence that this mechanism is not a simple enhancement of diffusive transport, but a fundamental transition in its character to intermittent, avalanche-like behavior. This transition is enabled by the collapse of the stabilizing shear, as shown previously in Fig. 5.
III.4.1 Statistical Signatures of Avalanche-like Transport,
The statistical properties of the core density fluctuations, measured by the RIP diagnostic at the midplane, reveal three distinct signatures of this transition.
First, the probability density function (PDF) of the fluctuation amplitudes evolves from Gaussian to heavy-tailed. Early in the discharge the PDF is Gaussian [yellow in Fig. 11(a)], characteristic of diffusive transport. As the density limit is approached, however, the PDF develops pronounced, non-Gaussian tails [purple in Fig. 11(a)]. This signifies a shift to a transport regime dominated by large-amplitude, intermittent events. The growing intermittency is quantified by the kurtosis of the PDF, which rises from the Gaussian value of 3 to about 6 [Fig. 11(b)].
@@ III.4.1 Statistical Signatures of Avalanche-like Transport,
Third, the transport acquires long-range temporal correlations, quantified by the Hurst exponent (), a measure of the persistence in a time series [Fig. 12(b)]. The Hurst exponent rises from , the value for uncorrelated, random-walk events (diffusive transport), to as the density limit is approached. A value of indicates persistent, correlated dynamics, another key feature of avalanche-like transport. This rise in is strongly correlated with a substantial decrease in the energy confinement factor , linking the emergence of this non-diffusive transport regime to the degradation in plasma performance.
Collectively, these three independent signatures—heavy-tailed PDFs, a power spectrum, and an elevated Hurst exponent—provide a consistent and compelling picture: core transport undergoes a fundamental transition from a diffusive to an intermittent, avalanche-like regime, which may be the mechanism for the core density limit.
@@ III.4.2 Core Density Saturation by Avalanche-like Turbulence
The emergence of avalanche-like dynamics provides a clear hypothesis for the core density limit: the large, intermittent transport events ultimately clamps the density profile. We test this hypothesis by correlating the core density (at ) with the RMS amplitude of the midplane turbulence measured by RIP.
@@
In the low-turbulence, high-confinement regime (, yellow points), which corresponds to the diffusive state, the core density rises robustly with increasing turbulence amplitude, following a strong power-law scaling of . This phase represents a state where fueling remains effective despite the presence of turbulence.
This behavior changes dramatically as the transport transitions to the high-turbulence, low-confinement regime (, purple points), which corresponds to the avalanche-like state. Here, the scaling weakens significantly, flattening to . This transition to a much weaker scaling demonstrates that the core fueling has become highly inefficient. The avalanche-like fluctuations effectively prevent any significant increases in density, leading to a saturation of the core density profile. This saturation occurs at densities far exceeding the Greenwald limit, approaching .
The weak overall scaling of density with turbulence () is consistent with the weak macroscopic power scaling of the core density limit () found previously. This alignment confirms that the core density limit is governed not by simple power balance, but by a fundamental transition in the turbulent transport itself, from diffusive to avalanche-like, which saturates the density profile.
@@ IV Conclusion and Discussion
This study re-examined the underlying physics of the density limit using negative triangularity plasmas on the DIII-D tokamak. Our findings reveal that the density limit, at least in NT plasmas without an H-mode edge, is not a singular disruptive boundary, but a multifaceted saturation phenomenon governed by distinct physical mechanisms: an edge density limit triggered by a radiative instability, and a core density limit driven by avalanche-like turbulent transport. Moreover, this study demonstrate sustained, non-disruptive operations at densities up to 1.8 times the Greenwald limit with substantial auxiliary heating.
@@
At the plasma edge, the density limit manifests as an abrupt saturation event triggered by a radiative instability on the high-field side. While geometrically similar to a disruptive MARFE, this radiative instability acts as a quasi-stable front that clamps the edge density at a level below the Greenwald limit, without terminating the discharge.
@@
In contrast, the core density is not directly constrained by the edge cooling and continues to rise well beyond the Greenwald limit. Its saturation is instead governed by a fundamental shift in the nature of core turbulence. As the density increases, core transport transitions from a diffusive state to one characterized by intermittent, avalanche-like dynamics. This is supported by signatures including heavy-tailed PDFs, -type power spectra, and elevated Hurst exponents. This persistent, non-diffusive transport regime prevents further increases in the core density.
@@
The dual mechanisms of regional density saturation are consistent with the disparate power scalings observed for the core () and edge (). The power-dependent nature of this limit and its link to transport physics show a compelling resemblance to density limits observed in stellarators. This suggests a potentially universal physics of power balance and transport that transcends specific magnetic configurations, pointing to a more fundamental basis for the density limit in magnetically confined plasmas.
@@
These findings also present a potential strategy for achieving high Greenwald fraction operation in future tokamaks, particularly those utilizing negative triangularity configurations. Success hinges on addressing the distinct physical mechanisms governing the density limit in the core and edge. The edge radiative instability can be mitigated through advanced divertor designs and impurity control aimed at managing boundary heat fluxes. Concurrently, advanced techniques to suppress core turbulence, such as strong rotational shear flows and Shafranov shift, can be developed to optimize density profiles and confinement. By pursuing these parallel control strategies, the operational space of future reactors can be extended into stable, high-density regimes well beyond conventional scaling.
@@ Acknowledgments
The authors gratefully acknowledge the invaluable support and contributions of the entire DIII-D team in performing this experiment. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards Nos. DE-FC02-04ER54698, DE-SC0019352, DE-SC0019004, DE-FG02-97ER54415, DE-AC52-07NA27344, and DE-SC0016154. The author (P.H.D.) acknowledges support from LLNL-led SciDAC ABOUND Project SCW1832. The author (O.S.) performed this work within the framework of the EUROfusion Consortium, via the Euratom Research and Training Programme (Grant Agreement No. 101052200– EUROfusion) and funded by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission, or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.

### v4

@@ III.3.3 Clamping of the Edge Density
The preceding observations suggest that the HFS radiation front can be a key trigger for the edge density saturation. To test this hypothesis quantitatively, we correlate the local edge density with the mode’s strength, again using the RMS amplitude of the line-integrated HFS density fluctuations () as a proxy.
Figure 9 plots the edge density (at ) against on a log-log scale. The data, compiled from multiple discharges, shows a bifurcation in the plasma behavior, defined by two distinct operational regimes that are distinguished by the energy confinement factor, .
@@
This linear relationship breaks down abruptly once the turbulence level surpasses a critical threshold, at which point the system transitions to a degraded-confinement state (low , purple points). While the turbulence amplitude continues to grow by nearly an order of magnitude, the edge density saturates, with its scaling flattening significantly ().
This bifurcation suggests that the edge density limit is not a gradual process, but a sharp transition associated with the onset of a radiative mode. Once initiated, the mode clamps the edge density at a value consistently below the Greenwald limit, leading to a significant increase in HFS turbulence. The apparent overall relationship across both regimes () also suggests the sharpness of this critical transition.
@@ III.3.4 Breakdown of Zonal Flow Regulation
The onset of the HFS radiation front has global consequences, leading to a failure of turbulence self-regulation across the plasma edge. We use BES measurements on the low-field side to observe this breakdown in detail (Fig. 10), revealing the microphysical mechanism responsible for enhanced transport. It is worth mentioning that BES measurements are taken from a different discharge (shot 194131) with lower NBI power, which results in a slightly earlier onset of the radiative cooling.
The trigger for this breakdown is the transition to the non-adiabatic regime. As the edge cools with the onset of the radiation front ( ms), the dimensionless adiabaticity parameter, , drops below unity [Fig. 10(a)]. This transition is expected to alter drift-wave turbulence, weakening the mechanisms that generate stabilizing zonal flows [40, 15].
@@
This breakdown in zonal flow regulation allows for increased particle loss. The outward turbulent particle flux [, Fig. 10(e)] increases significantly, while the turbulent spreading flux [, Fig. 10(f)] indicates that turbulence intensity is propagating inward from the separatrix. These measurements suggest a sequence: the formation of the HFS radiation front triggers edge cooling, which pushes the entire edge into a non-adiabatic state. This disables turbulence self-regulation and hence the zonal flow production on the LFS, enabling enhanced particle transport and contributing to the clamping of the edge density profile.
III.4 The Core Density Saturation and the Emergence of Avalanche-like Turbulence
While the edge density saturates below the Greenwald limit, the core density continues to rise but eventually plateaus at a value significantly above . This decoupling implies that a separate, core-localized mechanism must govern the core density profile. Here, we show that this mechanism is not a simple enhancement of diffusive transport, but is correlated with a transition in its character to intermittent, avalanche-like behavior. This transition can be enabled by the collapse of the stabilizing shear, as shown previously in Fig. 5.
III.4.1 Statistical Signatures of Avalanche-like Transport
The statistical properties of the density fluctuations, measured by the RIP diagnostic at the midplane, reveal three distinct signatures of this transition.
First, the probability density function (PDF) of the fluctuation amplitudes evolves from Gaussian to heavy-tailed. Early in the discharge the PDF is Gaussian [yellow in Fig. 11(a)], characteristic of diffusive transport. As the density limit is approached, the PDF develops non-Gaussian tails [purple in Fig. 11(a)]. This signifies a shift to a transport regime dominated by large-amplitude, intermittent events. The growing intermittency is quantified by the kurtosis of the PDF, which rises from the Gaussian value of 3 to about 6 [Fig. 11(b)].
@@ III.4.1 Statistical Signatures of Avalanche-like Transport
Third, the transport acquires long-range temporal correlations, quantified by the Hurst exponent (), a measure of the persistence in a time series [Fig. 12(b)]. The Hurst exponent rises from (random-walk events, diffusive transport) to as the density limit is approached. A value of indicates persistent, correlated dynamics, another key feature of avalanche-like transport. This increase in correlates with a substantial decrease in the energy confinement factor , linking the emergence of this non-diffusive transport regime to the degradation in plasma performance.
Collectively, these three independent signatures—heavy-tailed PDFs, a power spectrum, and an elevated Hurst exponent—are consistent with a scenario in which turbulent transport undergoes a transition from a diffusive to an intermittent, avalanche-like regime, which may be the mechanism for the core density limit.
It is worth noting that the RIP diagnostic is line-integrated, and its signal inherently contains contributions from the entire plasma profile along its midplane chord, including both the core and the edge. However, local measurements of the edge turbulence on the low-field side midplane from both DBS and BES do not exhibit the signatures of avalanche-like transport that we observe in the RIP data. Therefore, while the RIP signal contains contributions from the full profile, its change in character towards an avalanche-like state can serve as an indicator of the changing dynamics of the core transport.
@@ III.4.2 Core Density Saturation by Avalanche-like Turbulence
The emergence of avalanche-like dynamics provides a hypothesis for the core density limit: the large, intermittent transport events ultimately clamp the density profile. We test this hypothesis by correlating the core density (at ) with the RMS amplitude of the midplane turbulence measured by RIP.
@@
In the low-turbulence, high-confinement regime (, yellow points), which corresponds to the diffusive state, the core density rises with increasing turbulence amplitude, following a power-law scaling of . This phase represents a state where fueling remains effective despite the presence of turbulence.
This behavior changes as the transport transitions to the high-turbulence, low-confinement regime (, purple points), which corresponds to the avalanche-like state. Here, the scaling weakens significantly, flattening to . The avalanche-like fluctuations may effectively prevent any significant increases in density, leading to a saturation of the core density profile. This saturation occurs at densities far exceeding the Greenwald limit, approaching . The weak overall scaling of density with turbulence () is consistent with the weak macroscopic power scaling of the core density limit () found previously.
@@ IV Conclusion and Discussion
This study re-examined the underlying physics of the density limit using negative triangularity plasmas on the DIII-D tokamak. Our findings suggest that the density limit, at least in NT plasmas without an H-mode edge, is not a singular disruptive boundary, but a saturation phenomenon, or a ‘soft’ limit, governed by distinct physical mechanisms: an edge density limit triggered by radiative cooling, and a core density limit driven by avalanche-like turbulent transport. This study also demonstrates sustained, non-disruptive operations at densities up to 1.8 times the Greenwald limit with substantial auxiliary heating.
@@
At the plasma edge, the density limit manifests as an abrupt saturation event that is linked to the formation of a non-disruptive, radiation front on the high-field side. While geometrically similar to a disruptive MARFE, this radiative structure acts as a quasi-stable mode that leads to a non-adiabatic, high-turbulence regime and hence may clamp the edge density at a level below the Greenwald limit.
@@
In contrast, the core density is not directly constrained by the edge cooling and continues to rise beyond the Greenwald limit. Its saturation is instead associated with a fundamental shift in the nature of core turbulence. As the density increases, core transport transitions from a diffusive state to one characterized by intermittent, avalanche-like dynamics. This is supported by signatures including heavy-tailed PDFs, -type power spectra, and elevated Hurst exponents. This persistent, non-diffusive transport regime prevents further increases in the core density.
@@
The dual mechanisms of regional density saturation are consistent with the disparate power scalings observed for the core () and edge (). The power-dependent nature of this limit and its link to transport physics resembles density limits observed in stellarators. This suggests an underlying physics of power balance and transport that transcends specific magnetic configurations, pointing to a more fundamental basis for the density limit in magnetically confined plasmas.
@@
These findings also present a potential strategy for achieving high Greenwald fraction operation in future tokamaks, particularly those utilizing negative triangularity configurations. Success hinges on addressing the distinct physical mechanisms governing the density limit in the core and edge. The edge radiation front can be mitigated through advanced divertor designs and impurity control. Techniques to suppress core turbulence, such as strong rotational shear flows and Shafranov shift, can be developed to optimize density profiles and confinement. By using targeted controls for the distinct edge radiative and core turbulent limits, the operational space can be extended more efficiently than by relying on heating power alone.
A further question for the broader applicability of these findings is whether these power scaling and saturation mechanisms manifest in positive triangularity (PT) L-mode plasmas. However, conducting a comparable study in PT is challenging in present tokamaks. The exploration of a power-dependent density limit requires high auxiliary heating, which typically triggers a transition to H-mode in PT configurations. This moves the plasma into a different regime, where the H-mode density limit (HDL) is governed by distinct pedestal physics and not the L-mode mechanisms detailed in this work. Consequently, L-mode density limit studies in PT are restricted to low-power regimes that are often disruptive due to radiative condensation instabilities, preventing a clear investigation of the saturation phenomena. Further work is also required to establish the causal links between the HFS radiation front and the complex turbulence response.
@@ Acknowledgments
The authors gratefully acknowledge the invaluable support and contributions of the entire DIII-D team in performing this experiment. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards Nos. DE-FC02-04ER54698, DE-SC0019352, DE-SC0019004, DE-FG02-97ER54415, DE-AC52-07NA27344, DE-FG02-08ER54999 and DE-SC0016154. The author (P.H.D.) acknowledges support from LLNL-led SciDAC ABOUND Project SCW1832. The author (O.S.) performed this work within the framework of the EUROfusion Consortium, via the Euratom Research and Training Programme (Grant Agreement No. 101052200– EUROfusion) and funded by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission, or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.